OpenStudy (anonymous):

Solve the equation using square roots. x^2-9=0 A. –3, 3 B. –9, 9 C. D. no real number solutions

Parth (parthkohli):

\(\Rightarrow (x + 3)( x- 3) = 0\) Now use the zero-product rule. Do you know how to do so?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

no. how you do that?

Parth (parthkohli):

@pitachip Please don't answer directly, and welcome!

OpenStudy (anonymous):

x^2-9 = 0 x^2 = 9 x = +3, -3 Hope this helps.

OpenStudy (anonymous):

shut up

Parth (parthkohli):

\(\Rightarrow ( x + 3)( x - 3) = 0 \Longrightarrow x + 3 = 0, x - 3 = 0\)

Parth (parthkohli):

Solve both equations.

OpenStudy (anonymous):

I figured it out. thanks